Knight Inventory Systems, Inc., has announced a rights offer. The company has announced that it will take four rights to buy a new share in the offering at a subscription price of $29. At the close of business the day before the ex-rights day, the company’s stock sells for $50 per share. The next morning, you notice that the stock sells for $44 per share and the rights sell for $2 each.

What is the value of the stock ex-rights? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
What is the value of a right? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

  


Are the rights underpriced or overpriced?
  

  

What is the amount of immediate profit you can make on ex-rights day per share? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
Answer #1

a). PX = [NPRO + PS] / (N + 1)

= [(4 x $50) + $29] / (4 + 1) = $229 / 5 = $45.80

b). Value of a right = PRO - PX = $50 - $45.80 = $4.20

c). So, the rights are under-priced.

d). You can create an immediate profit on the ex-rights day if the stock is selling for $45.80 and the rights are selling for $2 by executing the following transactions:

Buy 4 rights in the market for 4($2) = $8. Use these rights to purchase a new share at the subscription price of $29. Immediately sell this share in the market for $45.80, creating an instant $8.80 profit.